# LSTP-Nav: Lightweight Spatiotemporal Policy for Map-free Multi-agent Navigation with LiDAR

### v4

@@ Abstract
Safe and efficient multi-agent navigation in dynamic environments remains inherently challenging, particularly when real-time decision-making is required on resource-constrained platforms. Ensuring collision-free trajectories while adapting to uncertainties without relying on pre-built maps further complicates real-world deployment. To address these challenges, we propose LSTP-Nav, a lightweight end-to-end policy for multi-agent navigation that enables map-free collision avoidance in complex environments by directly mapping raw 2D LiDAR point clouds to motion commands. At the core of this framework lies LSTP-Net (Lightweight Spatiotemporal Policy Network), an efficient network that processes raw LiDAR data using a gated recurrent unit-based architecture, enhanced with attention mechanisms to dynamically focus on critical environmental features while minimizing computational overhead. Additionally, a novel Heading Stability reward (HS reward) is introduced to optimize collision avoidance by incorporating angular velocity, prioritizing obstacles along the predicted heading, and enhancing training stability. To narrow the sim-to-real gap, we develop PhysReplay-Simlab, a multi-agent training simulator incorporating realistic physics, and a localized replay mechanism that actively mines near-failure experiences to improve learning efficacy. Relying solely on LiDAR data, LSTP-Nav achieves efficient zero-shot sim-to-real transfer on a CPU-only robotic platform, enabling robust navigation in dynamic and cluttered environments while maintaining computation frequencies above 40 Hz. Extensive experiments demonstrate that LSTP-Nav outperforms baselines with a 9.58% higher success rate and a 12.30% lower collision rate, underscoring its practicality and robustness for real-world applications. Project page: https://sites.google.com/view/xingrong2024efficient.
@@ II-B PhysReplay-Simlab
PhysReplay-Simlab integrates a local replay mechanism to address challenges in dense multi-agent scenarios, including spatiotemporal sparsity, emergent obstacle prediction, and rapid adaptation to dynamic obstacles. When an agent collides, it will be reset to its previous state from $N=300$ simulation steps, creating near-collision states. This accelerates collision data accumulation through focused adversarial sampling. To mitigate training instability caused by repeated short-term collisions, agents exceeding a threshold collision are randomly reset to valid states within an iteration, promoting exploration diversity while maintaining trajectory continuity.
@@ II-C2 Training Strategy
where $V_{targ}=A+V_{\phi}(s)$. The composite loss function is defined as $L^{{Total}}=L^{{P}}+\alpha L^{V}+\beta L^{E}$, where $\alpha$ and $\beta$ are empirically tuned hyperparameters. Network parameters are optimized using the Adam optimizer [32] with respect to $L^{Total}$. Finally, the algorithm terminates upon reaching the maximum iteration count in Line 23.
@@
where $w_{g}$ are the weights for $r_{g}$ and $d_{i}^{t}$ represents the minimum distance between agent $i$ and obstacles at timestep $t$. The goal reward provides dense incremental rewards for a progressive goal approach while reserving sparse, high-value rewards for final success.
@@
$r_{g}$ =$=\begin{cases}r_{collision}&\text{if }d_{i}^{t}<0.01\\ w_{c}\left(z_{max}-z_{i}^{t}\right)+r_{c}&\text{ otherwise}\end{cases},$ (16)
where $r_{arrival}=2$, $r_{collision}=-2$, $w_{g}=3.5$, $w^{\prime}_{c}=-0.1$, $w_{c}=-0.03$, $z_{max}$ denotes the maximum detection range of LiDAR, and $z_{i}^{t}$ represents the LiDAR observations for agent $i$ at timestep $t$. The weight vector $w_{c}$ follows a Gaussian distribution centered at the current angular displacement $\omega\Delta t$:
@@ III-A Experimental Setup
The proposed method (LSTP-Net) is evaluated through comparative experiments in both simulated and real-world environments, with the following baseline methods: NH-ORCA [33] and the CNN-based policy [26] with the conventional reward function. Ablation studies are performed to assess the contributions of key architectural components.
@@ III-B Simulation Experiments and Results
To comprehensively evaluate the LSTP-Net (ours), we conduct comparative experiments against established baselines: NH-ORCA [33] and CNN-based policy [26] with conventional rewards in simulated environments. Additionally, ablation studies are performed to isolate the contribution of key architectural components: (1) removing the attention mechanism (GRU policy), and (2) replacing the GRU with linear layers (Linear policy).

### v5

@@ Abstract
Safe and efficient multi-agent navigation in dynamic environments remains inherently challenging, particularly when real-time decision-making is required on resource-constrained platforms. Ensuring collision-free trajectories while adapting to uncertainties without relying on pre-built maps further complicates real-world deployment. To address these challenges, we propose LSTP-Nav, a lightweight end-to-end policy for multi-agent navigation that enables map-free collision avoidance in complex environments by directly mapping raw LiDAR point clouds to motion commands. At the core of this framework lies LSTP-Net, an efficient network that processes raw LiDAR data using a GRU architecture, enhanced with attention mechanisms to dynamically focus on critical environmental features while minimizing computational overhead. Additionally, a novel HS reward optimizes collision avoidance by incorporating angular velocity, prioritizing obstacles along the predicted heading, and enhancing training stability. To narrow the sim-to-real gap, we develop PhysReplay-Simlab, a physics-realistic multi-agent simulator, employs localized replay to mine near-failure experiences. Relying solely on LiDA, LSTP-Nav achieves efficient zero-shot sim-to-real transfer on a CPU-only robotic platform, enabling robust navigation in dynamic environments while maintaining computation frequencies above 40 Hz. Extensive experiments demonstrate that LSTP-Nav outperforms baselines with a 9.58% higher success rate and a 12.30% lower collision rate, underscoring its practicality and robustness for real-world applications.
@@ II-B PhysReplay-Simlab
PhysReplay-Simlab integrates a local replay mechanism to address challenges in dense multi-agent scenarios, including spatiotemporal sparsity, emergent obstacle prediction, and rapid adaptation to dynamic obstacles. As shown in Fig. 3, upon collision, the agent is reset to its state from N=300 simulation steps prior, generating near-collision scenarios for training. The replay mechanism accelerates collision data accumulation through focused adversarial sampling. To mitigate training instability caused by repeated short-term collisions, agents exceeding a threshold collision are randomly reset to valid states within an iteration, promoting exploration diversity while maintaining trajectory continuity.
@@ II-C2 Training Strategy
where $V_{\text{targ}}=A+V_{\phi}(s)$. The composite loss function is defined as $L^{{Total}}=L^{{P}}+\alpha L^{V}+\beta L^{E}$, where $\alpha$ and $\beta$ are empirically tuned hyperparameters. Network parameters are optimized using the Adam optimizer [32] with respect to $L^{Total}$. Finally, the algorithm terminates upon reaching the maximum iteration count in Line 23.
@@
where $w_{g}$ are the weight of $r_{g}$ and $d_{i}^{t}$ represents the minimum distance between agent $i$ and obstacles at timestep $t$. The goal reward provides dense incremental rewards for a progressive goal approach while reserving sparse, high-value rewards for final success.
@@
$r_{c}$ =$=\begin{cases}r_{collision}&\text{if }d_{i}^{t}<0.01\\ w_{c}\left(z_{max}-z_{i}^{t}\right)&\text{ otherwise}\end{cases},$ (16)
where $w_{c}$ are the weight of $r_{c}$, $z_{max}$ denotes the maximum detection range of LiDAR, and $z_{i}^{t}$ represents the LiDAR observations for agent $i$ at timestep $t$. The weight vector $w_{c}$ follows a Gaussian distribution centered at the current angular displacement $\omega\Delta t$:
@@ III-A Experimental Setup
The proposed method (LSTP-Net) is evaluated through comparative experiments in both simulated and real-world environments, with the following baseline methods: NH-ORCA [33] and the CNN policy [26] with the conventional reward function. Ablation studies are performed to assess the contributions of key architectural components.
@@ III-B Simulation Experiments and Results
To comprehensively evaluate the LSTP-Net (ours), we conduct comparative experiments against established baselines: NH-ORCA [33] and CNN policy [26] with conventional rewards in simulated environments. Additionally, ablation studies are performed to isolate the contribution of key architectural components: (1) removing the attention mechanism (GRU policy), and (2) replacing the GRU with linear layers (Linear policy).